OpenStudy (anonymous):

A 5.00-kg ball, moving to the right at a velocity of +2.00 m/s on a frictionless table, collides head-on with a stationary 7.50kg ball. find the final velocities of the ball if the collision is elastic

OpenStudy (irishboy123):

post the actual question 🏹

OpenStudy (anonymous):

@IrishBoy123 Perhaps you're having problems viewing the question? I can see it just fine from my side. Hmm, there might be an easier way, but this is probably how I would approach the question. Perhaps @IrishBoy123 can check it X) `Conservation of Momentum:`\[\huge Mv+m u=Mv_f+m u_f\]\[\huge 10+0=5v_f+7.5u_f\] `Conservation of Total Kinetic Energy:`\[\huge \frac{1}{2}Mv^2+\frac{1}{2}m u^2=\frac{1}{2}Mv_f^2+\frac{1}{2}m u_f^2\]\[\huge 10+0=2.5v_f^2+3.75u_f^2\] You should be able to solve for \(v_f\) and \(u_f\) as a system of two equations. \(v\) is the initial velocity of the first ball \(v_f\) is the final velocity of the first ball \(u\) is the initial velocity of the second ball \(u_f\) is the final velocity of the second ball \(M\) is the mass of the first ball \(m\) is the mass of the second ball You can also approximate that \(v_f \approx v_i\) and \(u_f \approx 2v_i\) and continue to solve from there. Source: http://hyperphysics.phy-astr.gsu.edu/hbase/colsta.html

OpenStudy (anonymous):

However, I think that approximation can only be made if the masses greatly different. 5kg and 7.5kg are kind of close to each other.